OpenStudy (anonymous):

Complete the equation of the graphed linear function. Write the slope in decimal form. y =[ ]x +[ ]

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

First, we can find the y-intercept, that's the place where it intersects on the y-axis, and what's that?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes, so we have: y = [_]x + 4

OpenStudy (anonymous):

Now for the slope, we can take any two points on the line and plug it into the slope formula.

OpenStudy (anonymous):

(0,4)?. . . or (5,0)

OpenStudy (anonymous):

We can use both. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{0-4}{5-0}\) Simplify that

OpenStudy (anonymous):

-4 and 5

OpenStudy (anonymous):

Yep, so the slope is -4/5 So our final answer is: y = -4/5x + 4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

oh and @iGreen. its a creeper from minecraft. . . if thats what you ment by "and what's that?"

OpenStudy (anonymous):

No I meant the y-intercept when I said that.

OpenStudy (anonymous):

oo ok. . . awkward